OpenStudy (anonymous):

help me plz guys every evening jenna empties her pockets and puts her change in a jar. at the end of the week she counts her money. one week she had 40 coins all of them dimes and quarters. when she added them up, she had a total of $7.75. let d = number of dimes and q =number of quarters. how many dimes did jenna have?

OpenStudy (anonymous):

25 quarters and 15 dimes

OpenStudy (anonymous):

so it would be 15 dimes

OpenStudy (anonymous):

what?! well thats a stupid question Jenna had 0 dimes because 7.75 can be made in quarters

OpenStudy (anonymous):

25 x .25 = 6.25 15 x .10 = 1.50 6.25 + 1.50 = 7.75

OpenStudy (anonymous):

Let Q = the number of quarters and D = the number of dimes. Then: 1) Q+D = 40 "One week she had 40 coins, all of them dimes and quarters...." 2) 0.25Q+0.1D = $7.75 "...she had a total of $7.75." Rewrite equation 1) as: D = 40-Q and substitute into equation 2). 2a) 0.25Q+0.1(40-Q) = 7.75 Simplify the left side. 2b) 0.25Q+4-0.1Q = $7.75 Subtract 4 from both sides. 2c) 0.25Q-0.1Q = 3.75 Combine the like-terms on the left side. 2d) 0.15Q = 7.75 Divide both sides by 0.15 2e) Q = 25 and: D = 40-Q D = 40-25 D = 15 Jenna had 25 quarters and 15 dimes. nvm http://www.algebra.com/algebra/homework/Problems-with-consecutive-odd-even-integers/Problems-with-consecutive-odd-even-integers.faq.question.204097.html

OpenStudy (anonymous):

I forgot to notice that there were 40 coins :p

OpenStudy (anonymous):

(:

OpenStudy (anonymous):

i will give you guys a another problem if you want

OpenStudy (anonymous):

sure why not

OpenStudy (anonymous):

unfortunately i'm going to be leaving soon, i have to go pick my cat up from the vet .

OpenStudy (anonymous):

alex stocks up for winter. he buys 24 cans of vegetables. he pays 80 cents per can for tomatoes and 40 cants per can for corn, for a total cost of $12.80. how many cans of corn does he buy?

OpenStudy (anonymous):

Let t = # cans tomatoes and c = # cans corn t + c = 24 subtract c from both sides c = 24 - t 80(t) + 40(c) = 1280 substitute in for c 80(t) + 40(24-t) = 1280 80t + 960 - 40t = 1280 subtract 960 from both sides 40t = 320 t = 8 c = 24 - t = 24 - 8 = 16 check work 80(8) + 40(16) = 1280 640 + 640 = 1280 1280 = 1280 He bought 16 cans of corn http://answers.yahoo.com/question/index?qid=20110524130003AAkBnoa

OpenStudy (anonymous):

ok well it was good having you here or I would have given the wrong answer lol @ilyautumn

OpenStudy (anonymous):

it is alright if i get the wrong answer

OpenStudy (anonymous):

haha no its not id feel bad lol

OpenStudy (anonymous):

i could aways take the test again

OpenStudy (anonymous):

haha id still feel bad if you would have to take the test again

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

this is my last question george and carmen went on a bicycle trip. they took a bus to their starting point, and then biked the rest. they travled 325 kilometers in total, and they biked kilometers more than they were bussed. let x = kilometers travled by bike and y = kilometers travled by bus. find how many kilomteers they travled by bike.